OpenStudy (anonymous):

Find the slope of the line containing each given pair of points. If the slope is undefined, state this. (−5, −1) and (2, −3)

OpenStudy (anonymous):

Slope = (Difference in y)/(Difference in x) or, \[Slope = (y_2 - y_1)/(x_2 - x_1)\] Can you solve it further or not??

OpenStudy (anonymous):

so (-5-2)/(-1-(-3)) ?

terenzreignz (terenzreignz):

The numerator should have the difference of the y-values, not the difference of the x-values... Check it again :)

OpenStudy (anonymous):

So how do I write it out I am confused

terenzreignz (terenzreignz):

Well your y-values are -1 and -3, and yet you put their difference in the denominator, how can you fix that? :)

OpenStudy (anonymous):

-1--3?

OpenStudy (anonymous):

You have given the points (x1, y1) and (x2, y2).. So, there y1 = -1 and y2 = -3 Can you find y2 - y1 = ????

terenzreignz (terenzreignz):

It's -3 -(-1) Can you do it from here? :)

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

y2 - y1 = -3 -(-1) = -3 + 1 = -2 Similarly find x2 - x1 = ????

OpenStudy (anonymous):

Yes you did it right y2 - y1 will be (-2)...

OpenStudy (anonymous):

-5- -1 or -5-1?

terenzreignz (terenzreignz):

yes of course now BE CAREFUL you selected -3 for your y2, so you HAVE TO select 2 as your x2... therefore x2 = 2 x1 = -5 so your x2 - x1 is... ? :)

OpenStudy (anonymous):

2-5= -3

OpenStudy (anonymous):

No, 2 - (-5) = ??? This will be the x2 - x1 i think...

OpenStudy (anonymous):

7

terenzreignz (terenzreignz):

You got it... so your slope is...?

OpenStudy (anonymous):

-2/7

OpenStudy (anonymous):

Yes you are correct dear...

OpenStudy (anonymous):

Thanks, so it is not an accurate slope?

terenzreignz (terenzreignz):

Ehh, we'll leave you to define what you mean by accurate slope :) It's the slope, there's really nothing more to it Although if you mean it isn't an integer value, then yes, it isn't "accurate" lol :D

OpenStudy (anonymous):

Okay thanks :)